OpenStudy (anonymous):

Hey everyone, I'm really confused on the concept of upper and lower bounds. An example would be: Suppose that f(3)=4 and 2(equalto or <) f'(x) (equalto or <) 3 for all x. Find an upper bound on the value of f(5). Please explain with steps so I can learn :) Thank you.

OpenStudy (anonymous):

ok we can thing of it like this. think that \[f'(x)\] is the speed limit. this says your speed is between 2 and 3. now at 3 pm you are at mile marker 4 on the highway. when can you be at 5 pm?

OpenStudy (anonymous):

5 pm is 2 hours later, so at most you could have gone \[2\times 3=6\] miles and now you are at mile marker 10 (since 6+4=10)

OpenStudy (anonymous):

or you could have traveled a minimum of \[2\times 2=4\] miles so if you went the minimum speed you would be at mile marker 8 (because 4+4=8)

OpenStudy (anonymous):

meaning \[8\leq f(5)\leq 10\]

OpenStudy (anonymous):

we can write this using the mean value theorem if you like.

OpenStudy (anonymous):

\[2\leq \frac{f(5)-f(3)}{5-3}\leq 3\] \[2\leq \frac{f(5)-4}{2}\leq 3\] \[4\leq f(5)-4\leq 6\] \[8\leq f(5)\leq 10\]

OpenStudy (anonymous):

Oh I get it, thanks a lot for you help!

OpenStudy (anonymous):

yw